OpenStudy (saeedd):

. Solid carbon is known to react with oxygen to produce carbon dioxide. If a mass of 1.2 g of carbon is burned in oxygen, 4.4 g of carbon dioxide is found to form. How many grams of oxygen gas reacted with the carbon?

OpenStudy (saeedd):

Ive been stuck on this for a while

OpenStudy (saeedd):

I seriously need help. I have a test tomorrow

OpenStudy (saeedd):

Well that seems really simple, and this is AP cHEM, SO IT SEEMS A BIT OFF.

OpenStudy (saeedd):

Sorry caps

OpenStudy (aveline):

Ah, so real chemistry. Have you gone over moles yet?

OpenStudy (australopithecus):

First off notice it is a 1 to 1 reaction. 1. Convert the product to moles moles = grams/molecular mass 2. Because it is 1 to 1 moles of CO2 produced = Moles of O2 reacted 3. Then use the formula: moles = grams/molecular mass to convert moles of O2 to grams of O2 Note its always a good idea to start from the products just in case there is not a complete reaction.

OpenStudy (australopithecus):

How to calculate molecular mass: https://www.youtube.com/watch?v=F9NkYSKJifs

OpenStudy (saeedd):

sooo

OpenStudy (aveline):

Molar Mass: Carbon: 12.01g/mol Oxygen (Diatomic): 16.00g/mol Carbon Dioxide: 44.01 g/mol C + O2 = CO2 We have 1.2 g of Carbon. Let's change it to moles: \[1.2gC \times \frac{ molC }{ 12.01gC } = 0.10molC\] We also have 4.4g of CO2 after the reaction. Let's also change this into moles. \[4.4gCO_2 \times \frac{ molCO_2 }{ 44.01gCO_2 } = 0.10molCO_2\]